# LNS2+RL: Combining Multi-agent Reinforcement Learning with Large Neighborhood Search in Multi-agent Path Finding

### v1

@@ Abstract
Multi-Agent Path Finding (MAPF) is a critical component of logistics and warehouse management, which focuses on planning collision-free paths for a team of robots in a known environment. Recent work introduced a novel MAPF approach, LNS2, which proposed to repair a quickly-obtainable set of infeasible paths via iterative re-planning, by relying on a fast, yet lower-quality, priority-based planner. At the same time, there has been a recent push for Multi-Agent Reinforcement Learning (MARL) based MAPF algorithms, which let agents learn decentralized policies that exhibit improved cooperation over such priority planning, although inevitably remaining slower. In this paper, we introduce a new MAPF algorithm, LNS2+RL, which combines the distinct yet complementary characteristics of LNS2 and MARL to effectively balance their individual limitations and get the best from both worlds. During early iterations, LNS2+RL relies on MARL for low-level re-planning, which we show eliminates collisions much more than a priority-based planner. There, our MARL-based planner allows agents to reason about past and future/predicted information to gradually learn cooperative decision-making through a finely designed curriculum learning. At later stages of planning, LNS2+RL adaptively switches to priority-based planning to quickly resolve the remaining collisions, naturally trading-off solution quality and computational efficiency. Our comprehensive experiments on challenging tasks across various team sizes, world sizes, and map structures consistently demonstrate the superior performance of LNS2+RL compared to many MAPF algorithms, including LNS2, LaCAM, and EECBS. In maps with complex structures, the advantages of LNS2+RL are particularly pronounced, with LNS2+RL achieving a success rate of over 50% in nearly half of the tested tasks, while that of LaCAM and EECBS falls to 0%. We finally experimentally validate our algorithm in a hybrid simulation of a warehouse mockup involving a team of 100 (real-world and simulated) robots.
1 Introduction
@@
Multi-Agent Path Finding (MAPF) focuses on the problem of generating a set of shortest collision-free paths for a team of agents from their initial positions to designated goals in a shared, known environment. It is NP-hard to solve optimally and is a core problem in many applications that may involve thousands of agents, such as warehouse automation and traffic management, where balancing solution quality and speed is crucial. Recently, Li et al. (2022) introduced LNS2, a large neighborhood search-based MAPF algorithm that iteratively reuse/fix currently infeasible paths via local re-planning, differing from mainstream MAPF algorithms that rely on more expensive search/backtracking. LNS2 exhibits greater scalability than systematic search planners Sharon et al. (2015) and higher solution quality than priority-based algorithms Erdmann & Lozano-Perez (1987). LNS2 first quickly generates an initial solution that may contain collisions, and then iteratively replans for carefully selected subsets of agents, using a fast, priority-based low-level planner (Safe Interval Path Planning with Soft Constraints (SIPPS) Li et al. (2022) for single agent pathfinding), to finally yield a collision-free solution. However, in high-agent-density tasks, the paths generated by this priority-based planner still often contain a large number of collisions, inevitably increasing the cumulative time needed to address the overall MAPF problem. In parallel to these developments, recent years have seen the development of many (slower) MAPF planners based on Multi-Agent Reinforcement Learning (MARL), which exhibit enhanced cooperation compared to priority-based planners. However, MARL-based planners inevitably exhibit performance degradation when deviating from their training setup (usually only 8 agents), which limits their performance on tasks containing a high density of agents. Moreover, the one-shot Success Rate (SR) of MARL-based planners often drops sharply due to a few agents failing to reach their goals within the time(step) limit, thereby wasting high-quality paths for the majority that succeeded.
@@
In this paper, we propose to combine large neighborhood search and MARL, and introduce a new algorithm called LNS2+RL. Our approach aims to balance the drawbacks of both LNS2 and MARL-based planners, showcasing enhanced performance across diverse scenarios. The overall framework of LNS2+RL follows that of LNS2, iteratively updating the overall solution with re-planned paths. However, it replaces the priority-based planner with a MARL-based planner during the challenging early stages of iterative re-planning to further improve solution quality (i.e., reduce collisions). Our MARL-based planner allows agents to access and reason about historical and predictive information, progressively mastering cooperative decision-making through a carefully crafted curriculum learning strategy. There, we rely on SIPPS to fix the paths planned by agents unable to reach their goals within a limited number of timesteps, thus ensuring the completeness of our MARL-based planner and maximizing the utilization of existing high-quality paths. LNS2+RL then adaptively switches to the original priority-based planner (according to the relative solution quality of the two algorithms in the previous re-planning step), to swiftly clean up remaining collisions, thus naturally trading-off solution quality and re-planning speed. Moreover, since the number of agents involved in each re-planning task (i.e., neighborhood size) is fixed to 8, LNS2+RL can naturally scale the overall planning to thousands of agents while keeping encountered tasks close to MARL’s training distribution.
Figure 1: Real portion of the warehouse with 10 robots (left), and complete rviz visualization with 100 robots (right). Real robots are circled in different colors (both in the real and simulated views), and crosses of the same colors represent their goal positions.
@@
We tested LNS2+RL on an extensive set of high-difficulty tasks with various team sizes, world sizes, and map structures. Our results indicate LNS2+RL consistently outperforms LNS2 in SR across nearly all tasks, maintaining a comparable Sum of Costs (SoC), and notably reducing the number of remaining Collision Pairs (CP) in the overall solution. Additionally, LNS2+RL surpasses EECBS Li et al. (2021) in SR across all tasks. Compared to LaCAM Okumura (2023), LNS2+RL exhibits a substantial advantage in SR on highly structured maps (achieving 100% SR compared to LaCAM’s 0%), while also achieving significantly lower SoC on less structured maps. Finally, as shown in Figure 1, we experimentally validated our planner in a hybrid simulation of a warehouse environment under real-world conditions, where we let a team of 100 agents (10 real-world, and 90 simulated ones) track the paths obtained by LNS2+RL, using an online execution framework Hönig et al. (2019).
2 Prior work
@@
MAPF algorithms developed in recent decades fall into three main categories based on solution optimality: optimal, bounded-suboptimal and unbounded-suboptimal. Common optimal algorithms include Conflict-Based Search (CBS) Sharon et al. (2015), joint-state A*, and their variants with additional speed-up techniques such as ICBS Boyarski et al. (2015), M* Wagner & Choset (2011), EPEA* Goldenberg et al. (2014), among others. Bounded-suboptimal algorithms trade off solution quality and runtime by finding a solution whose cost is at most $w$ times the optimal cost, where $w$ is a user-specified suboptimality factor. Such MAPF algorithms are usually variants of optimal MAPF algorithms, including EECBS Li et al. (2021), weighted EPEA* Barer et al. (2014), and inflated M* Wagner & Choset (2015), etc. Both optimal and bounded-suboptimal algorithms are able to ensure high solution quality but face scalability challenges when handling thousands of agents, often encountering timeouts or memory overflows. Conversely, unbounded-suboptimal algorithms (such as priority Erdmann & Lozano-Perez (1987) and rule-based planning Luna & Bekris (2011)) prioritize speed over optimality, scaling to thousands of agents easily, but may yield substantially suboptimal solutions in complex tasks and often do not guarantee completeness. More recently, learning-based approaches have begun to emerge. One class of approaches, exemplified by PRIMAL Sartoretti et al. (2019), directly learns a MAPF planner using MARL Ma et al. (2021); Wang et al. (2020), while another combines conventional algorithms with learning, exemplified by switching between two types of algorithms based on certain conditions Skrynnik et al. (2023) or by enhancing conventional algorithms’ components with learned strategies Huang et al. (2022).
3 Background
3.1 Multi-agent path finding
@@
In this paper, we consider a common formulation of the MAPF problem, defined within a 2d 4-edge grid $G=(V,E)$, encompassing a set of $m$ agents $A=\{a_{1},\ldots,a_{m}\}$ and a finite set of static obstacles. Each agent $a_{i}$ initiates from a distinct starting position $s_{i}\in V$, with the objective of reaching its designated goal $g_{i}\in V$. At each discretized timestep, an agent can either move to an adjacent vertex or wait at its current vertex. The path $p_{i}$ for each agent is a sequence of adjacent or identical vertices $(v_{1}^{i},\ldots,v_{k}^{i})$ satisfying $v_{1}^{i}=s_{i}$, $v_{k}^{i}=g_{i}$. We assume that the agents wait at their goal vertices until all agents have reached their goals. A solution of MAPF $P=\left\{p_{i}\mid a_{i}\in A\right\}$ is a path set for all agents, ensuring all agents reach their goals at a certain timestep without any vertex (agents occupy the same vertex) or swapping collisions (agents traverse the same edge simultaneously). This paper evaluates solution quality using a sum of costs metric, defined as the sum of timesteps required by each agent to reach its goal $\sum_{1\leq i\leq m}\left|p_{i}\right|$.
3.2 LNS2
@@
LNS2 is an unbounded-suboptimal and memory-efficient MAPF algorithm based on large neighborhood search. It begins by planning a set of paths that may contain collisions, iteratively selecting subsets for re-planning through an Adaptive Large Neighborhood Selection (ALNS) method based on previous experience. A replanning solver then computes new paths for these selected agents and evaluates whether to adopt the new paths depending on whether they reduce or maintain the number of CP within the overall solution. This replanning process continues iteratively until the overall path set is collision-free or the time limit is exceeded.
@@
The low-level, re-planning solver employed in LNS2 is a variant of priority planning (termed PP+SIPPS), which utilizes SIPPS algorithm to plan individual paths for agents in a random priority order. SIPPS aims to find a path for an agent to avoid collisions with hard obstacles (static obstacles) while minimizing collisions with soft obstacles (pre-existing paths). It runs five times faster than Space-Time A*, ensuring completeness and returning the shortest path with zero soft collisions if one exists. However, in dense-obstacle scenarios, SIPPS may produce a path with significantly more soft collisions than the minimum due to its heuristic function ignoring soft collisions that occur when the agent waits within a safe interval that contains soft obstacles. Despite the fast speed of SIPPS, this reduced solution quality necessitates a higher number of iterations to fully solve high-difficulty MAPF instances, resulting in substantial cumulative time expenditure. Therefore, there is an urgent need to develop a new re-planning solver that can generate high-quality solutions (fewer collisions) while potentially sacrificing speed, to find a better balance between quality and time throughout the overall planning.
4 Re-planning tasks as a MARL problem
4.1 Path finding with mixed dynamic obstacles
@@
The re-planning task can be defined as the Path Finding with Mixed Dynamic Obstacles (PMDO) problem. In the PMDO problem, two finite sets of obstacles are considered. The first category, hard obstacles, aligns with the immovable obstacles in conventional MAPF. The second category, soft obstacles, refers to other unselected agents and is characterized by the ability to move to any vertex not occupied by hard obstacles and remain there for an arbitrary duration. Notably, the complete trajectory of each soft obstacle is fully disclosed and remains constant throughout the re-planning task. Any occupied vertex at a given timestep $t$ can contain either a single hard obstacle or multiple soft obstacles/agents. The objective of the re-planning task is to find the shortest path for each selected agent from their start vertex to the goal vertex while avoiding collisions with hard obstacles and minimizing collisions with both soft obstacles and other agents.
4.2 RL environment setup
@@
Aligned with previously outlined MAPF and PMDO problems, we consider a discrete 2D 4-connected gridworld environment where the size of each agent, goal, and obstacle is one grid cell. Each agent can take a discrete action at every timestep, which includes moving one unit towards any of the four cardinal directions or staying still. The agents’ actions are executed synchronously, with collision checks conducted subsequently to the execution of all actions. Actions resulting in collisions with hard obstacles or exceeding world boundaries are classified as invalid. Only valid actions are sampled during training and evaluation, and an additional supervised loss is added to the overall MARL loss.

### v2

@@ Abstract
Multi-Agent Path Finding (MAPF) is a critical component of logistics and warehouse management, which focuses on planning collision-free paths for a team of robots in a known environment. Recent work introduced a novel MAPF approach, LNS2, which proposed to repair a quickly obtained set of infeasible paths via iterative replanning, by relying on a fast, yet lower-quality, prioritized planning (PP) algorithm. At the same time, there has been a recent push for Multi-Agent Reinforcement Learning (MARL) based MAPF algorithms, which exhibit improved cooperation over such PP algorithms, although inevitably remaining slower. In this paper, we introduce a new MAPF algorithm, LNS2+RL, which combines the distinct yet complementary characteristics of LNS2 and MARL to effectively balance their individual limitations and get the best from both worlds. During early iterations, LNS2+RL relies on MARL for low-level replanning, which we show eliminates collisions much more than a PP algorithm. There, our MARL-based planner allows agents to reason about past and future information to gradually learn cooperative decision-making through a finely designed curriculum learning. At later stages of planning, LNS2+RL adaptively switches to PP algorithm to quickly resolve the remaining collisions, naturally trading off solution quality (number of collisions in the solution) and computational efficiency. Our comprehensive experiments on high-agent-density tasks across various team sizes, world sizes, and map structures consistently demonstrate the superior performance of LNS2+RL compared to many MAPF algorithms, including LNS2, LaCAM, EECBS, and SCRIMP. In maps with complex structures, the advantages of LNS2+RL are particularly pronounced, with LNS2+RL achieving a success rate of over 50% in nearly half of the tested tasks, while that of LaCAM, EECBS and SCRIMP falls to 0%.
Codehttps://github.com/marmotlab/LNS2-RL
Introduction
@@
Multi-Agent Path Finding (MAPF) focuses on the problem of generating a set of shortest collision-free paths for a team of agents from their initial positions to designated goals in a shared environment. It is a core problem in many applications, such as warehouse automation, where balancing solution quality and speed is crucial. A recent paper (Li et al. 2022) introduced LNS2, a large neighborhood search-based MAPF algorithm that iteratively replans currently infeasible paths, differing from mainstream MAPF algorithms that rely on more expensive search/backtracking. LNS2 exhibits greater scalability than systematic search algorithms (Sharon et al. 2015) and prioritized planning algorithms (Erdmann and Lozano-Perez 1987). It first quickly generates an initial solution that may contain collisions, and then iteratively replans for carefully selected subsets of agents, using a fast, low-level Prioritized Planning algorithm based on Safe Interval Path Planning with Soft Constraints (termed PP+SIPPS) to finally yield a collision-free solution. However, in challenging tasks, the paths generated by PP+SIPPS still often contain a large number of collisions, inevitably increasing the cumulative time needed to address the overall MAPF problem.
In parallel to these developments, recent years have seen the development of many (slower) MAPF planners based on Multi-Agent Reinforcement Learning (MARL), which exhibit enhanced cooperation compared to prioritized planning algorithms. However, MARL-based planners inevitably exhibit performance degradation when deviating from their training setup (usually only 8 agents). Moreover, the one-shot Success Rate (SR) of MARL-based planners often drops sharply due to a few agents failing to reach their goals within the timestep limit, thereby wasting high-quality paths for the majority that succeeded.
@@
In this paper, we propose combining large neighborhood search and MARL, introducing a new algorithm called LNS2+RL, which balances the drawbacks of both LNS2 and MARL-based planners. The overall framework of LNS2+RL follows that of LNS2, iteratively updating the overall solution with re-planned paths. However, during the challenging early stages of iterative replanning, it replaces PP+SIPPS with a MARL-based planner, generating paths with fewer collisions at a slower speed. There, we rely on SIPPS to fix the paths planned by agents unable to reach their goals within a limited number of timesteps, thus ensuring the completeness of our MARL-based planner and maximizing the utilization of existing high-quality paths. LNS2+RL then adaptively switches to the original PP+SIPPS to swiftly clean up remaining collisions, naturally trading-off solution quality (number of collisions in the solution) and replanning speed. Moreover, since the number of agents involved in each replanning task (i.e., neighborhood size) is fixed to 8, LNS2+RL can naturally scale the overall planning to thousands of agents while keeping encountered tasks close to MARL’s training distribution.
@@
We trained only a single MARL model but tested LNS2+RL on an extensive set of high-agent-density tasks, where the team sizes and map structures differed from those used during training. Our results indicate that LNS2+RL consistently outperforms LNS2 in SR across nearly all tasks, maintaining a comparable Sum of Costs (SoC), and notably reducing the number of remaining Colliding Pairs (CP) in the overall solution. Additionally, LNS2+RL surpasses the search-based bounded-suboptimal algorithm EECBS (Li, Ruml, and Koenig 2021) and MARL-based planner SCRIMP (Wang et al. 2023) in SR across all tasks. Compared to the search-based unbounded-suboptimal algorithm LaCAM (Okumura 2023), LNS2+RL exhibits a substantial advantage in SR on maps with complex obstacle structure (achieving 100% SR compared to LaCAM’s 0%), while also achieving significantly lower SoC on maps with simple obstacle structures. Finally, we experimentally validated our planner in a hybrid simulation of a warehouse environment under real-world conditions.
Prior Work
@@
MAPF algorithms developed in recent decades fall into three main categories based on solution optimality: optimal, bounded-suboptimal, and unbounded-suboptimal. Common optimal algorithms include Conflict-Based Search (CBS) (Sharon et al. 2015), joint-state A*, and their variants with additional speed-up techniques such as ICBS (Boyarski et al. 2015), M* (Wagner and Choset 2011), EPEA* (Goldenberg et al. 2014), among others. Bounded-suboptimal algorithms trade off solution quality and runtime by finding a solution whose cost is at most $w$ times the optimal cost, where $w$ is a user-specified suboptimality factor. Such MAPF algorithms are usually variants of optimal MAPF algorithms, including EECBS (Li, Ruml, and Koenig 2021), and inflated M* (Wagner and Choset 2015), etc. Both optimal and bounded-suboptimal algorithms are able to ensure high solution quality but face scalability challenges when handling thousands of agents, often encountering timeouts or memory overflows. Conversely, unbounded-suboptimal algorithms (such as prioritized planning (Erdmann and Lozano-Perez 1987) and rule-based algorithm (Luna and Bekris 2011)) prioritize speed over optimality, scaling to thousands of agents easily, but may yield substantially suboptimal solutions in complex tasks and often do not guarantee completeness.
More recently, learning-involved approaches have begun to emerge. One class of approaches, exemplified by PRIMAL (Sartoretti et al. 2019), leverages MARL or imitation learning to directly train MAPF planners. These approaches often incorporate non-learning algorithms to further improve the performance of the learning-based planners by enriching observations (Damani et al. 2021), designing rewards (Skrynnik et al. 2024), or generating training data (Li et al. 2021). Although these learning-based planners emphasize their superior scalability compared to optimal or bounded-suboptimal search-based algorithms, nearly all of them still fail to outperform scalable suboptimal non-learning algorithms like LNS2 within the same time constraints. Another class of methods seeks to enhance conventional algorithms by using machine learning techniques to replace manually crafted heuristics or components, exemplified by learning a priority ordering for prioritized planning (Zhang et al. 2022), selecting subsets of agents that need to be replanned (Huang et al. 2022). However, these approaches do not directly optimize the path planner itself.
Background
Multi-Agent Path Finding
@@
In this paper, we consider a common formulation of the MAPF problem, defined within a 2D 4-neighbor grid with a finite set of static obstacles, and encompassing a set of $m$ agents $A=\{a_{1},\ldots,a_{m}\}$. Each agent $a_{i}$ initiates from a distinct start $s_{i}$, with the objective of reaching its designated goal $g_{i}$. At each discretized timestep, an agent can either move to an adjacent cell not occupied by an obstacle or wait at its current cell. The path $p_{i}$ with length $k$ (where the path length $k$ may vary for different paths) for each agent is a sequence of adjacent or identical cells $(c_{1}^{i},\ldots,c_{k}^{i})$ satisfying $c_{1}^{i}=s_{i}$, $c_{k}^{i}=g_{i}$. We assume that the agents wait at their goals until all agents have reached their goals. A solution of MAPF $P=\left\{p_{i}\mid a_{i}\in A\right\}$ is a path set for all agents, ensuring all agents reach their goals without any vertex (agents occupy the same cell at the same timestep) or swapping collisions (agents simultaneously attempting to move into each other’s current cells). This paper adopts the sum of costs metric as one of the evaluation criteria, which is defined as the sum of timesteps required by each agent to reach its goal $\sum_{1\leq i\leq m}\left|p_{i}\right|$.
LNS2
@@
LNS2 (Li et al. 2022) is an unbounded-suboptimal and memory-efficient MAPF algorithm based on large neighborhood search. It begins by planning a set of paths that may contain collisions, and then iteratively selects subsets for replanning using adaptive neighborhood selection based on past experience. A replanning solver then computes new paths for these selected agents and evaluates whether to adopt the new paths depending on whether they reduce or maintain the number of CP within the overall solution. This replanning process continues iteratively until the overall path set is collision-free or the time limit is exceeded.
@@
The low-level replanning solver employed in LNS2 is a prioritized planning algorithm PP+SIPPS, which utilizes SIPPS to sequentially plan paths for each agent based on a randomly assigned priority order, where lower-priority agents need to avoid the paths already planned for higher-priority agents as much as possible. SIPPS is designed to find a path for an agent that avoids collisions with hard obstacles (static obstacles) while minimizing collisions with soft obstacles (pre-existing paths). It ensures completeness and returns the shortest path with zero soft collisions if one exists. However, in dense-obstacle scenarios, SIPPS may produce a path with significantly more soft collisions than the minimum due to its heuristic function ignoring soft collisions that occur when the agent waits within a safe interval (a contiguous period of time) that contains soft obstacles. More importantly, the prioritized planning algorithm relies on a static-priority-based conflict avoidance mechanism, which significantly reduces the flexibility of the planning process. This limitation results in suboptimal paths and a larger number of collisions for lower-priority agents, making it difficult to achieve global cooperation. Despite the fast speed of PP+SIPPS, this reduced solution quality (more collisions) necessitates a higher number of iterations to fully solve high-difficulty MAPF instances, resulting in substantial cumulative time expenditure. Therefore, there is an urgent need to develop a new replanning solver that can generate high-quality solutions (fewer collisions) while potentially sacrificing speed, to find a better balance between quality and time throughout the overall planning.
Path Finding with Mixed Dynamic Obstacles
@@
The replanning task of LNS2 for selected agents can be defined as the Path Finding with Mixed Dynamic Obstacles (PMDO) problem. In the PMDO problem, two finite sets of obstacles are considered. The first category, hard obstacles, aligns with the immovable obstacles in conventional MAPF. The second category, soft obstacles, refers to other unselected agents and is characterized by the ability to move to any cell not occupied by hard obstacles and remain there for an arbitrary duration. Notably, the complete path of each soft obstacle is fully disclosed and remains constant throughout the replanning task. Any occupied cell at a given timestep $t$ can contain either a single hard obstacle or multiple soft obstacles/agents. The objective of the replanning task is to find a path for each selected agent from its start to its goal that avoids collisions with hard obstacles, prioritizes minimizing collisions with soft obstacles and other selected agents, and then minimizes the path length.
Algorithm 1 LNS2+RL
1:  Input: A MAPF instance $I$
2:  $P=\left\{p_{i}\mid a_{i}\in A\right\}\leftarrow$ lnitialSolver $(I)$
3:  $Q=\emptyset$
4:  $t_{m}=\max\{|p_{i}|:p_{i}\in P\}$, $t_{l}=d_{l}\cdot t_{m}$, $t_{h}=d_{h}\cdot t_{m}$
5:  while runtime limit not exceeded and $CP(P)>0$ do
6:     $A_{s},P^{-}\leftarrow\operatorname{selectAgentSet}(I,P)$
7:     if $|Q|<\mu$ or $\bar{Q}>\rho$ then
8:        $P^{rl},\hat{P}\leftarrow\operatorname{MARLPlanner}\left(I,P\setminus P^{-},A_{s},t_{l}\right)$
9:        $A_{r}\leftarrow\{a_{i}\in A_{s}:p_{i,\text{end}}^{rl}\neq g_{i}\}$
10:        $P^{rl}\leftarrow\operatorname{PrioritizedPlanner}\left(I,P\setminus P^{-},P^{rl},A_{r},t_{l}\right)$
11:        $A_{h}\leftarrow\{a_{i}\in A_{s}:|p_{i}^{rl}|\geq t_{h}\}$
12:        $P^{rl}\leftarrow\operatorname{PrioritizedPlanner}\left(I,P\setminus P^{-},P^{rl},A_{h},0\right)$
13:        Update $Q$ according to if $CP(P^{rl})$geq $CP(\hat{P})$
14:        if $CP(\left(P\backslash P^{-}\right)\cup P^{rl})\leq CP(P)$ then
15:           $P\leftarrow\left(P\backslash P^{-}\right)\cup P^{rl}$
16:           continue
17:        end if
18:     else
19:        $\hat{P}\leftarrow\operatorname{PrioritizedPlanner}\left(I,P\setminus P^{-},\emptyset,A_{s},0\right)$
20:     end if
21:     if $CP(\left(P\backslash P^{-}\right)\cup\hat{P})\leq CP(P)$ then
22:        $P\leftarrow\left(P\backslash P^{-}\right)\cup\hat{P}$
23:     end if
24:  end while
25:  return $P$
Method: LNS2+RL
In this section, we first present the overall framework of LNS2+RL and then detail the MARL model training.
Overall Framework
The strengths and weaknesses of PP+SIPPS and our MARL-based planner are distinct yet complementary. PP+SIPPS is fast and effective for simple tasks but tends to generate paths with more soft collisions in complex tasks. Conversely, our MARL-based planner, although slower (approximately seven times slower), excels at planning paths with fewer collisions in difficult tasks (about one-fourth of PP+SIPPS, with detailed data provided in “Performance analysis“ subsection). Therefore, LNS2+RL algorithm retains the overarching framework of LNS2 but uses the MARL-based planner in the initial, more challenging stages of iterative planning for better solution quality (fewer collisions). It then adaptively switches to PP+SIPPS to quickly clean up the remaining collisions, thus balancing solution quality (number of collisions in the solution) with the speed of the replanning task solvers. Moreover, the one-shot SR of MARL-based planners often drops sharply because a few agents fail to reach their goals within the constrained episode length, even though most agents have already found high-quality paths. To address this, we employ PP+SIPPS to supplement paths for these few agents, ensuring the completeness of the final solution.
The pseudocode for the LNS2+RL algorithm is presented in Algorithm 1. Specifically, for solving a MAPF instance $I$, LNS2+RL starts by invoking the initial solver (PP+SIPPS) to generate a path set $P$ containing collisions (lines 2). Furthermore, LNS2+RL introduces a queue $Q$ with a maximum capacity of $\mu$, used to store the comparison results of the number of CP in the paths generated by PP+SIPPS and the MARL planner, enabling adaptive switching (line 3). Subsequently, LNS2+RL counts the maximum length of individual paths in the initial solution ($t_{m}$), sets the timestep threshold $t_{l}$ for pausing the MARL-planner, along with the maximum allowable path length $t_{h}$, where the constant value $d_{h}>d_{l}$ (line 4). If the termination criteria are not met, LNS2+RL iteratively selects a subset of agents $A_{s}\subseteq A$ using adaptive neighborhood selection (the same as in LNS2) and removes their paths $P^{-}=\left\{p_{i}\in P:a_{i}\in A_{s}\right\}$ from $P$, treating the remaining part of the path set $P\backslash P^{-}$ as fixed to construct a PMDO replanning task (lines 6). The neighborhood size (the size of $A_{s}$) is fixed at 8 (equivalent to the number of agents used in MARL training). If the size of $Q$ is below $\mu$, or its average $\bar{Q}$ exceeds another threshold $\rho$, LNS2+RL infers that the benefit of using the MARL-planner exceeds that of PP+SIPPS at this stage (line 7). Therefore, LNS2+RL chooses the trained MARL planner as the replanning solver. The function in line 8 uses the MARL-planner to generate actions until timestep $t_{l}$ (or until the replanning task is fully resolved, i.e., all agents have reached their goal without future soft collisions). It yields two path sets: $\hat{P}$, derived by PP+SIPPS at the onset of the replanning task, acting as the reference paths for the MARL policy, and path set $P^{rl}$, computed by the MARL policy. For incomplete paths in $P^{rl}$ where agents haven’t reached their goals (Line 9), we employ PP+SIPPS to plan the remaining paths from $t_{l}$ and merge these new paths with the existing ones (Line 10). Subsequently, all path lengths in $P^{rl}$ are reviewed, and those exceeding $t_{h}$ are directly replaced with paths re-planned by PP+SIPPS from timestep zero (lines 11-12). This operation aims to remove non-contributory segments planned by the MARL planner, thus reducing the SoC of the final solution $P$ and planning time for subsequent iterations. Next, LNS2+RL updates $Q$ by appending 0 if $CP(P^{rl})\geq CP(\hat{P})$, and 1 otherwise (lines 13). Finally, LNS2+RL first checks whether $P^{rl}$ meets the update condition (line 14), namely if the CP of the repaired path set $\left(P\backslash P^{-}\right)\cup P^{rl}$ is no larger than that of the old path set. If not, LNS2+RL continues to check whether $\hat{P}$ meets the update condition (line 21). The MARL planner continues to re-plan until either the switching condition or termination condition is met. Once the replanning solver switches to PP+SIPPS, it cannot switch back to the MARL-planner and always addresses remaining collisions using the same procedure as LNS2.
Replanning Tasks as a MARL Problem
RL Environment Setup
@@
Aligned with previously outlined MAPF and PMDO problems, we consider a discrete 2D 4-neighbor grid environment where the size of each agent, goal, and obstacle is one grid cell. The agents’ actions that result in collisions with hard obstacles or exceed world boundaries are classified as invalid. Only valid actions are sampled during training and evaluation, and an additional supervised loss is added to the overall MARL loss to reduce the probability of choosing invalid actions.